# Poisoning Prevention in Federated Learning and Differential Privacy via Stateful Proofs of Execution

### v2

@@ I Introduction
As illustrated in Fig. 1, a crucial security obstacle in the adoption of these schemes is an adversary ($\sf{\mathcal{A}dv}$) that feeds back-end aggregators with forged data, sabotaging the entire collection process even when the other participants (other edge devices and the back-end) are honest. These attacks, known as poisoning attacks, have been widely recognized and are currently considered a major problem for both FL [53] and LDP [7]. Existing mitigations are either data-driven (e.g., in the case of data poisoning detection [57, 23, 6]) or algorithmic, e.g., by making FL and LDP mechanisms more resilient against these attacks [50, 32, 55, 33]. In both cases, security is by design best-effort, since these approaches cannot detect/prevent data poisoning at its source, i.e., at the edge devices themselves (typically, resource-constrained IoT devices).
@@
Verifier ($\sf{\mathcal{V}rf}$) Prover ($\sf{\mathcal{P}rv}$) 1) Request to run $\sf{\mathcal{F}}$ 2) ${\sf{\mathcal{O}}}\leftarrow{\sf{\mathcal{F}}}()$ 3) Compute $\sigma$ 4) Respond with ($\sigma$, $\sf{\mathcal{O}}$) 5) Verify($\sigma$, $\sf{\mathcal{O}}$)
(a) $\sf{PoX}$ protocol steps
Verifier ($\sf{\mathcal{V}rf}$) Prover ($\sf{\mathcal{P}rv}$) 1) Request to run ($\sf{\mathcal{F}}$$\sf{\mathcal{I}}$) 2) ${\sf{\mathcal{O}}}\leftarrow{\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ 3) Compute $\sigma$ 4) Respond with ($\sigma$, $\sf{\mathcal{O}}$) 5) Verify($\sigma$, $\sf{\mathcal{O}}$) no $\sf{\mathcal{S}}$ leakage
(b) $\sf{PoSX}$ protocol steps
Figure 2: Standard $\sf{PoX}$ vs $\sf{PoSX}$
@@ I-A On the Insufficiency of Classic $\sf{PoX}$ to Avert Poisoning
As shown in Fig. 3, PoX is a challenge-response protocol composed of the following steps:
@@ I-B Our Contributions
To address the aforementioned limitations, this work introduces the security notion of Proof of Stateful Execution ($\sf{PoSX}$) to enable input validation and state preservation, in addition to classic PoX guarantees. The former relaxes the constraint of inputless functions in traditional $\sf{PoX}$, while the latter ensures that $\sf{PoSX}$ can use $\sf{\mathcal{P}rv}$ pre-existing states as long as they originate from a prior authentic $\sf{PoSX}$ execution. In essence, PoSX offers assurance that execution of $\sf{\mathcal{F}}$, computed with authentic input $\sf{\mathcal{I}}$ and state $\sf{\mathcal{S}}$, denoted ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$, occurred faithfully, without disclosing $\sf{\mathcal{S}}$ to $\sf{\mathcal{V}rf}$. Figure 1(b) highlights key differences between $\sf{PoX}$ and $\sf{PoSX}$ at a high level.
@@
We show that $\sf{SLAPP}$ can support a wide range of data collection schemes, including FL and LDP, all with poisoning-free guarantees. Compared to prior data-centric mitigations in FL [50, 6, 33] or LDP [7], $\sf{SLAPP}$ offers two key benefits: it is agnostic to the underlying collection scheme (and implementation thereof), thereby capable of supporting both FL and LDP without changes in its trusted code base (TCB); it approaches the problem from a system security perspective, requiring no assumptions about the underlying data and its distribution. In summary, our anticipated contributions are the following:
@@
    Real-World Prototypes: To validate $\sf{SLAPP}$ and foster reproducibility, we provide three implementations, each utilizing distinct cryptographic schemes: symmetric, traditional asymmetric, and quantum-resistant asymmetric primitives. These implementations are prototyped and opensourced [3] using a real-world IoT development board: SAML11 Xplained Pro [36].
@@
    Evaluation: We conduct various experiments to assess $\sf{SLAPP}$’s efficiency. Our results demonstrate small runtime and memory overhead atop the baseline, in exchange for increased security and flexibility. Finally, we provide detailed case studies highlighting $\sf{SLAPP}$’s efficacy in thwarting poisoning attacks within FL/LDP.
@@ II-A Privacy-Preserving Data Collection Schemes
Local Differential Privacy-based Data Collection (LDP-DC): The goal of this scheme is for an IoT sensor, $\sf{\mathcal{P}rv}$, to output a noisy sensor value ${\sf{\mathcal{O}}}$ such that it preserves $\epsilon$-LDP; informally speaking, $\epsilon$-LDP guarantees that ${\sf{\mathcal{O}}}$ leaks no information about its original value ${\sf{\mathcal{O}}}^{\prime}$ except with a small probability constrained by $\epsilon$. At the same time, collecting ${\sf{\mathcal{O}}}$ from a large number of $\sf{\mathcal{P}rv}$-s allows $\sf{\mathcal{V}rf}$ to obtain certain statistics of the collected sensor values with high confidence. Various LDP mechanisms have been proposed to achieve $\epsilon$-LDP in different settings. As an example case, we focus on the LDP mechanism called Basic RAPPOR [22] noting that the concepts introduced in this work can be generalized to other LDP-based mechanisms.
@@ 1 Func ${\sf{\mathcal{F}}}$, ${\sf{\mathcal{I}}}$ $f$, $p$, $q$):
2       ${\sf{\mathcal{O}}}^{\prime}\leftarrow{\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$ $b\leftarrow UE({\sf{\mathcal{O}}}^{\prime})$ $b^{\prime}\leftarrow PRR(b,f,B)$ ${\sf{\mathcal{O}}}\leftarrow IRR(b^{\prime},p,q)$ return $\sf{\mathcal{O}}$
3
@@ 1 Func $\sf{\mathcal{F}}$, $\sf{\mathcal{I}}$):
2       ${\sf{\mathcal{O}}}\leftarrow{\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$ $D$.append($\sf{\mathcal{O}}$)
3
@@ 6             $W\leftarrow W-\alpha\cdot\nabla(W;D)$
8      ${\sf{\mathcal{O}}}\leftarrow W$ return $\sf{\mathcal{O}}$
9
@@ 8      ${\sf{\mathcal{O}}}\leftarrow W$ return $\sf{\mathcal{O}}$
Federated Learning-based data collection (FL-DC): Contrary to LDP-DC, FL-DC [30] requires $\sf{\mathcal{P}rv}$ to send a locally trained machine learning (ML) partial model to $\sf{\mathcal{V}rf}$ instead of sensor readings. This keeps raw sensor data local to $\sf{\mathcal{P}rv}$ and not directly accessible by $\sf{\mathcal{V}rf}$.
A typical FL-DC consists of two phases depicted in Algorithm 2. In the first phase, $\sf{\mathcal{P}rv}$ invokes the Sense-Store function to collect a raw sensor reading and store it in a local list $D$. This phase can be repeated multiple times over a certain period to collect more training data on $\sf{\mathcal{P}rv}$. Once a sufficient amount of training data is gathered, $\sf{\mathcal{P}rv}$ receives globally trained weights $W$ from $\sf{\mathcal{V}rf}$ along with the training parameters ($t$ and $\alpha$). It then triggers the Train function that utilizes $W$ as a base model to train the data in $D$. This function outputs the locally trained weights ${\sf{\mathcal{O}}}$ without disclosing $D$ to $\sf{\mathcal{V}rf}$.
@@ II-B ARM TrustZone-M Security Extensions
  • Hardware-enforces World Isolation. TrustZone-M ensures complete isolation of these worlds by implementing several hardware controls (i.e., TZASC, TZMA modules) to enforce access control to hardware resources (e.g., program and data memory, peripherals) for these two worlds. With isolation in place, TrustZone ensures the Non-Secure World is unable to access any code and data located in the Secure World. This assures that the Secure world remains secure even if an adversary can fully modify or compromise the Non-Secure world software state.
@@ III-A Network and Usage Model
We consider an IoT setting as shown earlier in Figure 1, consisting of two entity types: one $\sf{\mathcal{V}rf}$ and multiple $\sf{\mathcal{P}rv}$-s. $\sf{\mathcal{P}rv}$ is a resource-constrained sensor deployed in a physical space of interest, e.g., a smart home, office, or factory. $\sf{\mathcal{V}rf}$ is a remote service provider that orchestrates these $\sf{\mathcal{P}rv}$-s. As an example, $\sf{\mathcal{P}rv}$-s could be smart light bulbs that are used in many smart homes and can be controlled by the end-user through Philips $\sf{\mathcal{V}rf}$’s application services; or Samsung could act as $\sf{\mathcal{V}rf}$ that provides a service for the end-user to command all SmartThings-compatible devices.
Besides offering this service, $\sf{\mathcal{V}rf}$ wishes to collect sensor data generated by $\sf{\mathcal{P}rv}$-s to further improve the service performance or enhance the user experience. In the context of this work, $\sf{\mathcal{V}rf}$ has the option to employ LDP-DC or FL-DC as its preferred data collection scheme, depending on the desired outcome and privacy considerations. As poisoning attacks could sabotage the collection outcome, $\sf{\mathcal{V}rf}$ also aims to detect and prevent such attacks to safeguard the authenticity of the outcomes.
@@ III-B Adversary Model
Finally, in line with the standard LDP and FL $\sf{\mathcal{A}dv}$ models, we also consider the possibility of malicious $\sf{\mathcal{V}rf}$. In the latter, $\sf{\mathcal{A}dv}$’s goal is to learn sensitive data on $\sf{\mathcal{P}rv}$, which could be accomplished by exploiting weaknesses in the underlying data collection scheme.
@@ IV $\sf{PoSX}$ and Associated Definitions
[PoX Security]Let $\sf{\mathcal{F}}$ represent an arbitrary software function (code) execution of which is requested by $\sf{\mathcal{V}rf}$ on $\sf{\mathcal{P}rv}$, producing output ${\sf{\mathcal{O}}}$. A protocol is considered PoX-secure if and only if the protocol outputting $\top$ implies:
@@
L1 – Lack of input validation. Definition IV considers inputless $\sf{\mathcal{F}}$ functions. This is because classic $\sf{PoX}$ does not support verification of the origin and integrity of inputs received by $\sf{\mathcal{P}rv}$. This limitation is significant for applications where $\sf{\mathcal{V}rf}$ must provide input $\sf{\mathcal{I}}$ as part of $\sf{\mathcal{F}}$ execution on $\sf{\mathcal{P}rv}$ (e.g., FL/LDP cases discussed in Sec. II-A).
@@
Remark: note that when $\sf{\mathcal{S}}$ contains only public information, L2 can be obviated by L1 by making $\sf{\mathcal{S}}$ a part of the authenticated output of a $\sf{PoX}$ instance and used as a $\sf{\mathcal{V}rf}$-defined input to $\sf{\mathcal{F}}$ in a subsequent $\sf{PoX}$ instance. However, when $\sf{\mathcal{S}}$ contains information that must remain hidden from $\sf{\mathcal{V}rf}$ (the case of our target applications – recall Sec. II-A), consistency of $\sf{\mathcal{S}}$ must be ensured locally at $\sf{\mathcal{P}rv}$, making L1 and L2 independent challenges.
@@ V-A Overview of $\sf{SLAPP}$ Workflow
Building on TrustZone-M hardware-enforced world isolation (recall Sec. II-B), our approach is to implement $\sf{SLAPP}$’s RoT in and execute it from the Secure World. Meanwhile, normal applications are untrusted (hereby referred to as untrusted software) and reside in the Non-Secure World.
@@
$\sf{SLAPP}$ workflow is depicted in Figure 3. At a high level, it enforces the following operation sequence upon receiving a $\sf{PoSX}$ request from $\sf{\mathcal{V}rf}$.:
@@
  • In order to access $\sf{\mathcal{S}}$ in ❻, ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ must call CheckState function in the Secure World. CheckState takes one input argument, representing the current $\sf{\mathcal{S}}$ value. Its task is to authenticate this value by matching it against the latest benign $\sf{\mathcal{S}}$ value from a prior execution stored within the Secure World.
@@ V-B $\sf{SLAPP}$ in Detail
Protocol V-B details the $\sf{\mathcal{V}rf}$leftrightarrow $\sf{\mathcal{P}rv}$ interation in $\sf{SLAPP}$
@@
An instance of $\sf{SLAPP}$ protocol starts when $\sf{\mathcal{V}rf}$ generates a $\sf{PoSX}$ request, in Step 1, comprising: (i) an identifier ${\sf{\mathcal{F}}}$ for which function to execute; (ii) input arguments $\sf{\mathcal{I}}$; and (iii) a monotonically-increasing counter-based challenge $c_{\sf{\mathcal{V}rf}}$. This request instructs $\sf{\mathcal{P}rv}$ to execute ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ in the Non-Secure World on $\sf{\mathcal{P}rv}$ and return the result to $\sf{\mathcal{V}rf}$ along with proof of this expected execution. Following this, $\sf{\mathcal{V}rf}$ signs this request (via Sign function), producing token $\sigma_{\sf{\mathcal{V}rf}}$, and attaches this token to the request before sending it to $\sf{\mathcal{P}rv}$ in Step 2 of Protocol V-B.
@@
If all checks succeed, Execute updates the local counter with $c$ with $c_{\sf{\mathcal{V}rf}}$ and initializes three Secure World flags in Step 5:

### v3

# SLAPP: Poisoning Prevention in Federated Learning and Differential Privacy via Stateful Proofs of Execution

@@ I Introduction
As illustrated in Fig. 1, a crucial security obstacle in the adoption of these schemes is an adversary ($\sf{\mathcal{A}dv}$) that feeds back-end aggregators with forged data, sabotaging the entire collection process even when the other participants (other edge devices and the back-end) are honest. These attacks, known as poisoning attacks, have been widely recognized and are currently considered a major problem for both FL [56] and LDP [6]. Existing mitigations are either data-driven (e.g., in the case of data poisoning detection [6, 5, 33, 22]) or algorithmic, e.g., by making FL and LDP mechanisms more resilient against these attacks [3, 24, 23, 28, 52]. In both cases, security is by design best-effort, since these approaches cannot detect/prevent data poisoning at its source, i.e., at the edge devices themselves (typically, resource-constrained IoT devices).
@@ I-A On the Insufficiency of Classic $\sf{PoX}$ to Avert Poisoning
PoX is a challenge-response protocol composed of the following steps:
@@ I-B Our Contributions
To address the aforementioned limitations, this work introduces the security notion of Proof of Stateful Execution ($\sf{PoSX}$) to enable input validation and state preservation, in addition to classic PoX guarantees. The former relaxes the constraint of inputless functions in traditional $\sf{PoX}$, while the latter ensures that $\sf{PoSX}$ can use $\sf{\mathcal{P}rv}$ pre-existing states as long as they originate from a prior authentic $\sf{PoSX}$ execution. In essence, PoSX offers assurance that execution of $\sf{\mathcal{F}}$, computed with authentic input $\sf{\mathcal{I}}$ and state $\sf{\mathcal{S}}$, denoted ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$, occurred faithfully, without disclosing $\sf{\mathcal{S}}$ to $\sf{\mathcal{V}rf}$.
@@
We show that $\sf{SLAPP}$ can support a wide range of data collection schemes, including FL and LDP, all with poisoning-free guarantees. Compared to prior data-centric mitigations in FL [52, 5, 34] or LDP [6], $\sf{SLAPP}$ offers two key benefits: it is agnostic to the underlying collection scheme (and implementation thereof), thereby capable of supporting both FL and LDP without changes in its trusted code base (TCB); it approaches the problem from a system security perspective, requiring no assumptions about the underlying data distributions. In summary, our anticipated contributions are the following:
@@
    Real-World Prototypes: To validate $\sf{SLAPP}$ and foster reproducibility, we provide three implementation variants, each utilizing distinct cryptographic schemes: symmetric, traditional asymmetric, and quantum-resistant primitives. These implementations are prototyped and opensourced [47] on a real-world IoT development board: NUCLEO-L552ZE-Q [51].
@@
    Evaluation: We conduct various experiments to assess $\sf{SLAPP}$’s efficiency. Our results demonstrate small runtime and memory overhead atop the baseline, in exchange for increased security and flexibility. Finally, we provide detailed case studies highlighting $\sf{SLAPP}$’s efficiency and efficacy in thwarting poisoning attacks within FL/LDP.
@@ II-A Privacy-Preserving Data Collection Schemes
Local Differential Privacy-based Data Collection (LDP-DC). The goal of this scheme is for an IoT sensor, $\sf{\mathcal{P}rv}$, to output a noisy sensor value ${\sf{\mathcal{O}}}$ such that it preserves $\epsilon$-LDP; informally speaking, $\epsilon$-LDP guarantees that ${\sf{\mathcal{O}}}$ leaks no information about its original value ${\sf{\mathcal{O}}}^{\prime}$ except with a small probability constrained by $\epsilon$. At the same time, collecting ${\sf{\mathcal{O}}}$ from a number of $\sf{\mathcal{P}rv}$-s allows $\sf{\mathcal{V}rf}$ to obtain certain statistics of the collected sensor values with high confidence. Various LDP mechanisms have been proposed to achieve $\epsilon$-LDP in different settings. As an example case, we focus on the LDP mechanism called Basic RAPPOR [21] noting that the concepts introduced in this work can be generalized to other LDP-based mechanisms.
@@ 2 Func ${\sf{\mathcal{F}}}$, ${\sf{\mathcal{I}}}$ $f$, $p$, $q$):
3       ${\sf{\mathcal{O}}}^{\prime}\leftarrow{\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$
4       $b\leftarrow UE({\sf{\mathcal{O}}}^{\prime})$
5       $b^{\prime}\leftarrow PRR(b,f,B)$
6       ${\sf{\mathcal{O}}}\leftarrow IRR(b^{\prime},p,q)$
7       return $\sf{\mathcal{O}}$
8
9
@@ 2 Func $\sf{\mathcal{F}}$, $\sf{\mathcal{I}}$):
3       ${\sf{\mathcal{O}}}\leftarrow{\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$
4       $D$.append($\sf{\mathcal{O}}$)
5
6
7
@@ 10             $W\leftarrow W-\alpha\cdot\nabla(W;D)$
13      ${\sf{\mathcal{O}}}\leftarrow W$
14       return $\sf{\mathcal{O}}$
15
16
17
@@ 13      ${\sf{\mathcal{O}}}\leftarrow W$
Federated Learning-based data collection (FL-DC). Contrary to LDP-DC, FL-DC [31] requires $\sf{\mathcal{P}rv}$ to send a locally trained machine learning (ML) partial model to $\sf{\mathcal{V}rf}$ instead of sensor readings. This keeps raw sensor data local to $\sf{\mathcal{P}rv}$ and not directly accessible by $\sf{\mathcal{V}rf}$.
A typical FL-DC consists of two phases depicted in Algorithm 2. In the first phase, $\sf{\mathcal{P}rv}$ invokes the Sense-Store function to collect a raw sensor reading and store it in a local list $D$. This phase can be repeated multiple times over a certain period to collect more training data on $\sf{\mathcal{P}rv}$. Once sufficient training data is gathered, $\sf{\mathcal{P}rv}$ receives globally trained weights $W$ from $\sf{\mathcal{V}rf}$ along with the training parameters ($t$ and $\alpha$). It then triggers the Train function that utilizes $W$ as a base model to train the data in $D$. This function outputs the locally trained weights ${\sf{\mathcal{O}}}$ without disclosing $D$ to $\sf{\mathcal{V}rf}$.
@@ II-B ARM TrustZone-M Security Extensions
  • Hardware-enforces World Isolation. TrustZone-M ensures complete isolation of these worlds by implementing several hardware controls (i.e., TZASC/TZMA) to enforce access control to hardware resources (e.g., program and data memory, peripherals) for these two worlds. With isolation in place, TrustZone ensures the Non-Secure World is unable to access any code and data located in the Secure World. This assures that the Secure world remains secure even if an adversary can fully modify or compromise the Non-Secure world software state.
@@ III-A Network and Usage Model
We consider an IoT setting as shown earlier in Fig. 1, consisting of two entity types: one $\sf{\mathcal{V}rf}$ and multiple $\sf{\mathcal{P}rv}$-s. $\sf{\mathcal{P}rv}$ is a resource-constrained sensor deployed in a physical space of interest, e.g., a smart home, office, or factory. $\sf{\mathcal{V}rf}$ is a remote service provider that orchestrates these $\sf{\mathcal{P}rv}$-s. As an example, $\sf{\mathcal{P}rv}$-s could be smart light bulbs that are used in many smart homes and can be controlled by the end-user through Philips $\sf{\mathcal{V}rf}$’s application services; or Samsung could act as $\sf{\mathcal{V}rf}$ that provides a service for the end-user to command all SmartThings-compatible devices.
Besides offering this service, $\sf{\mathcal{V}rf}$ wishes to collect sensor data generated by $\sf{\mathcal{P}rv}$-s to further improve the service performance or enhance the user experience. In this work, $\sf{\mathcal{V}rf}$ has the option to employ LDP-DC or FL-DC as its preferred data collection scheme, depending on the desired outcome and privacy considerations. As poisoning attacks could sabotage the collection outcome, $\sf{\mathcal{V}rf}$ also aims to detect and prevent such attacks to safeguard the authenticity of the outcomes.
@@ III-B Adversary Model
Further, we consider $\sf{\mathcal{A}dv}$ to be adaptive [57], i.e., it is aware of the algorithm and all the specifics of the protocol executed between $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$. As a result, $\sf{\mathcal{A}dv}$ is allowed to modify its attack strategy by modifying $\sf{\mathcal{P}rv}$’s software state (except for hardware-enforced protections) and the communication between $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ to attempt to circumvent the proposed defense.
Finally, in line with the standard LDP and FL $\sf{\mathcal{A}dv}$ models, we also consider the possibility of malicious $\sf{\mathcal{V}rf}$. In the latter, $\sf{\mathcal{A}dv}$’s goal is to learn sensitive data on $\sf{\mathcal{P}rv}$ while executing the protocol.
@@ IV $\sf{PoSX}$ and Associated Definitions
[PoX Security [11]]Let $\sf{\mathcal{F}}$ represent an arbitrary software function (code) execution of which is requested by $\sf{\mathcal{V}rf}$ on $\sf{\mathcal{P}rv}$, producing output ${\sf{\mathcal{O}}}$. A protocol is considered PoX-secure if and only if the protocol outputting $\top$ implies:
@@
L1 – Lack of input validation. Definition IV considers inputless $\sf{\mathcal{F}}$ functions. This is because classic $\sf{PoX}$ does not support verification of the origin and integrity of inputs received by $\sf{\mathcal{P}rv}$. This limitation is significant for applications where $\sf{\mathcal{V}rf}$ must provide input $\sf{\mathcal{I}}$ as part of $\sf{\mathcal{F}}$ execution on $\sf{\mathcal{P}rv}$ (e.g., FL/LDP cases discussed in Section II-A).
@@
Remark: when $\sf{\mathcal{S}}$ contains only public information, L2 can be obviated by L1 by making $\sf{\mathcal{S}}$ a part of the authenticated output of a $\sf{PoX}$ instance and used as a $\sf{\mathcal{V}rf}$-defined input to $\sf{\mathcal{F}}$ in a subsequent $\sf{PoX}$ instance. However, when $\sf{\mathcal{S}}$ must remain hidden from $\sf{\mathcal{V}rf}$ (the case of our target applications – recall Section II-A), consistency of $\sf{\mathcal{S}}$ must be ensured locally at $\sf{\mathcal{P}rv}$, making L1 and L2 independent challenges.
@@ V-A Overview of $\sf{SLAPP}$ Workflow
Building on TrustZone-M hardware-enforced world isolation (recall Section II-B), our approach is to implement $\sf{SLAPP}$’s RoT in and execute it from the Secure World. Meanwhile, normal applications are untrusted (hereby referred to as untrusted software) and reside in the Non-Secure World.
@@
$\sf{SLAPP}$ workflow is depicted in Fig. 2. At a high level, it enforces the following operation sequence upon receiving a $\sf{PoSX}$ request from $\sf{\mathcal{V}rf}$.:
@@
  • Before accessing $\sf{\mathcal{S}}$ in ❻, ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ must call CheckState in the Secure World. CheckState takes one input argument, representing the current $\sf{\mathcal{S}}$ value. Its task is to authenticate this value by matching it against the latest benign $\sf{\mathcal{S}}$ value from a prior execution stored within the Secure World.
@@ V-B $\sf{SLAPP}$ in Detail
Protocol V-B details the $\sf{\mathcal{V}rf}$leftrightarrow $\sf{\mathcal{P}rv}$ interation in $\sf{SLAPP}$.
@@
An instance of $\sf{SLAPP}$ protocol starts when $\sf{\mathcal{V}rf}$ generates a $\sf{PoSX}$ request, in Step 1, comprising: (i) an identifier ${\sf{\mathcal{F}}}$ for which function to execute; (ii) input arguments $\sf{\mathcal{I}}$; and (iii) a monotonically-increasing counter-based challenge $c_{\sf{\mathcal{V}rf}}$. This request instructs $\sf{\mathcal{P}rv}$ to execute ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ in the Non-Secure World on $\sf{\mathcal{P}rv}$ and return the result to $\sf{\mathcal{V}rf}$ along with proof of this expected execution. Following this, $\sf{\mathcal{V}rf}$ signs this request (via Sign function), producing token $\sigma_{\sf{\mathcal{V}rf}}$, and attaches this token to the request before sending it to $\sf{\mathcal{P}rv}$ in Step 2.
@@
If all checks succeed, Execute updates the local counter with $c_{\sf{\mathcal{V}rf}}$ and initializes three Secure World flags in Step 5:
@@
Protocol 1: $\sf{SLAPP}$ Protocol
Verifier ($\sf{\mathcal{V}rf}$) (1) Generate an authenticated $\sf{PoSX}$ request: $\sigma_{\sf{\mathcal{V}rf}}\leftarrow\textsf{Sign}({\sf{sk_{{\sf{\mathcal{V}rf}}}}},\textsf{H}({\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}}))$ (2) Send ($\sigma_{\sf{\mathcal{V}rf}}$, $\sf{\mathcal{F}}$, $\sf{\mathcal{I}}$, $c_{\sf{\mathcal{V}rf}}$) to $\sf{\mathcal{P}rv}$ Prover ($\sf{\mathcal{P}rv}$) In Non-Secure World: (3) Call Execute in Secure World with the received request In Execute function, Secure World: (4) Authenticate the request and abort if $r=\bot$ or $exec=\top$: $r\leftarrow(c_{\sf{\mathcal{V}rf}}>c)\land\textsf{Verify}({\sf{pk_{{\sf{\mathcal{V}rf}}}}},\sigma_{\sf{\mathcal{V}rf}},{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}})$ (5) Update counter: $c\leftarrow c_{\sf{\mathcal{V}rf}}$ and initialize: $exec\leftarrow\top$, $stateUsed\leftarrow\bot$ $stateChecked\leftarrow\bot$ (6) Disable interrupts and measure $PMEM$ and $\sf{\mathcal{V}rf}$ request: $h\leftarrow\textsf{H}(PMEM,{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}})$ (7) Call ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ in Non-Secure World In $\sf{\mathcal{F}}$ function, Non-Secure World: (8) Run ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$. Pass control to the Secure World when CheckState($s$) is called In CheckState function, Secure World: (9) Perform an integrity check on $s$, store the result to $p$ and return to Non-Secure World: $stateChecked\leftarrow(\textsf{H}(s)\stackrel{{\scriptstyle?}}{{=}}{\sf{\mathcal{S}}}_{sec})$. Also, set $stateUsed\leftarrow\top$. In $\sf{\mathcal{F}}$ function, Non-Secure World: (10) Continue with ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ execution. Pass control to the Secure World when SetState($s$) is called In SetState function, Secure World: (11) Securely set ${\sf{\mathcal{S}}}_{sec}$ based on input $s$ and return to Non-Secure World: ${\sf{\mathcal{S}}}_{sec}\leftarrow\textsf{H}(s)\text{ if }(stateChecked\land exec)$ In $\sf{\mathcal{F}}$ function, Non-Secure World: (12) Continue with ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ until the execution is completed, producing output $\sf{\mathcal{O}}$, and then return to its caller, Execute, with $\sf{\mathcal{O}}$ In Execute function, Secure World: (13) Abort if $(exec~{}\land stateUsed~{}\land\neg stateChecked)$. Otherwise, include $\sf{\mathcal{O}}$ to the measurement and compute the proof: $\sigma\leftarrow\textsf{Sign}({\sf{sk_{{\sf{\mathcal{P}rv}}}}},\textsf{H}(h,{\sf{\mathcal{O}}}))$ (14) Reset all flags: $exec\leftarrow\bot,stateChecked\leftarrow\bot,stateUsed\leftarrow\bot$ (15) Enable interrupts and return with $({\sf{\mathcal{O}}},\sigma)$ In Non-Secure World: (16) Forward $({\sf{\mathcal{O}}},\sigma)$ to $\sf{\mathcal{V}rf}$ Verifier ($\sf{\mathcal{V}rf}$) (17) Increment $c_{\sf{\mathcal{V}rf}}$ and validate $\sigma$ by: $r\leftarrow\textsf{ValidatePoSX}({\sf{pk_{{\sf{\mathcal{P}rv}}}}},\sigma,PMEM^{\prime},{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}},{\sf{\mathcal{O}}})$ The protocol outputs $r$.